OpenStudy (anonymous):

Solve for x, where x is a real number 4e^2x=12 could any1 please explain how to solve this step by step. I really want to be able to understand this. Is the answer even a real number?

OpenStudy (solomonzelman):

\(\large\color{blue}{4e^{2x}=12 \LARGE\color{white}{ \rm │ }}\) \(\large\color{blue}{e^{2x}=3 \LARGE\color{white}{ \rm │ }}\) \(\large\color{blue}{\ln(e^{2x})=\ln(3) \LARGE\color{white}{ \rm │ }}\) \(\large\color{blue}{2x\ln(e)=\ln(3) \LARGE\color{white}{ \rm │ }}\) ( `ln(e)=1` ) \(\large\color{blue}{2x=\ln(3) \LARGE\color{white}{ \rm │ }}\) \(\large\color{blue}{x=\ln(3)\div 2 \LARGE\color{white}{ \rm │ }}\)

OpenStudy (mrnood):

divide by 4 e^2x = 3 take loges (base e) 2x = ln3 you got it?...

OpenStudy (anonymous):

thanks guys, i got it now :)

OpenStudy (solomonzelman):

You can play around with \(\large\color{blue}{x=\ln(3)\div 2 \LARGE\color{white}{ \rm │ }}\) \(\large\color{blue}{x=\ln(3)\div 2 \LARGE\color{white}{ \rm │ }}\) \(\large\color{blue}{x=\ln(3)\div 2\ln(e) \LARGE\color{white}{ \rm │ }}\) \(\large\color{blue}{x=\ln(3)\div \ln(e^2) \LARGE\color{white}{ \rm │ }}\) \(\large\color{blue}{x=\ln(3-e^2) \LARGE\color{white}{ \rm │ }}\)